OpenStudy (anonymous):

If (x-2A)(2x+1) = 2x^2 - Bx + 10 A * B = ?

OpenStudy (polaris_s0i):

\[(x-2A)(2x+1) = 2x^2 - Bx + 10 \] \[2x^2 + x -4Ax -2A = 2x^2 - Bx + 10 \] \[2x^2 + (1-4A)x -2A = 2x^2 - Bx + 10 \] \[\cancel{2x^2} + (1-4A)x -2A = \cancel{2x^2} - Bx + 10 \] you can now see that: \[-2A = 10\] \[A = -\frac{10}{2}\] \[A=-5\] substitute into the equation for B: \[B = 1-4A\] \[B = 1-4(-5)\] \[B = 20\]